A box contains 5 white balls, 3 black balls, and 2 red balls.
A-What is the probability of drawing a white ball?
B- How many white balls must be added to the box so that the probability of drawing a white ball is 3/4?
C-How many black balls must be added to the ORIGINAL assortment so that the probability of drawing a white ball is 1/4?
The answer for A is .5, and the answer for B is 10; however, what is the answer for C! I need to show my work...Please help me with this q... Thank you

Answer by stanbon(75887)

A box contains 5 white balls, 3 black balls, and 2 red balls.
A-What is the probability of drawing a white ball?
B- How many white balls must be added to the box so that the probability of drawing a white ball is 3/4?
C-How many black balls must be added to the ORIGINAL assortment so that the probability of drawing a white ball is 1/4?
The answer for A is .5, and the answer for B is 10; however, what is the answer for C! I need to show my work...Please help me with this q...
===============
Let that number of added black balls be "x":
Solve for "x":
(5)/(10+x) = 1/4
---
10 + x = 20
x = 10 (# of additional black balls needed)
================
Cheers,
Stan H.